### v4

[…]
Generalized Hypercube Queueing Models with Overlapping Service Regions
[…]
We present a generalized hypercube queueing model, building upon the original model by Larson (1974), focusing on its application to overlapping service regions, such as police beats. To design a service region, we need to capture the workload and police car operation, a type of mobile server. The traditional hypercube queueing model excels in capturing the dynamics of systems with light traffic, as it primarily considers whether each server is busy or idle. However, contemporary service operations often experience saturation, in which each server in the system can only process a subset of calls, and a queue in front of each server is allowed. Hence, the simple binary status for each server becomes inadequate, prompting the need for a more intricate state space analysis. Our proposed model addresses this challenge by utilizing a Markov model with a large state space, represented by non-negative integer-valued vectors, along with a truncated hyperlattice queueing model approximation. By exploiting the sparsity of the transition matrix, we efficiently compute the steady-state distribution of the truncated hyperlattice model and show that it provides a close approximation under canonical dispatching policies. This solution can then be used to evaluate general performance metrics for the service system. We validate the effectiveness of our model through simulations of various artificial service systems. We also apply our model to the Atlanta police operational system, which faces challenges such as an increased workload, significant staff shortages, and the impact of boundary effects on crime incidents. Using real 911 calls-for-service data, our analysis indicates that a police operations system with permitted overlapping patrols can significantly mitigate these problems, leading to more effective deployment of the police force. Although the paper focuses on police districting applications, the generalized hypercube queueing model is applicable to other mobile server models in the general setup.
[…]
Hypercube queueing model; Overlapping service regions
1 Introduction
Service systems, such as Emergency Service Systems (ESS), are vital in providing emergency aid to communities, protecting public health, and ensuring safety. However, these systems often face resource limitations, including constrained budgets and personnel shortages (Peng et al. 2020, Yoon et al. 2021, Zhu et al. 2022). As a result, the public and associated organizations need to evaluate and improve the efficiency of these systems to maintain their effectiveness.
Service systems are generally designed such that a specific emergency response unit is primarily responsible for a small geographical area while remaining available to aid neighboring regions if necessary, either through a structured dispatch hierarchy or on an ad-hoc basis. For instance, in policing service systems, personnel are typically assigned to a designated area known as a beat. Police departments often allocate patrol forces by dividing a city’s geographical areas into multiple police zones. The design of these patrol zones has a significant impact on response times to 911 calls. Officers patrol their assigned beats and respond to emergency calls, with each spatial region having one queue.
A key component for analyzing service systems is to model the queueing dynamics of mobile servers, such as police cars, ambulances, and shared rides, depending on the context. Mobile servers refer to resources or servers that are not fixed in one location but are mobile, often to provide services in varying locations or under changing conditions. Some examples of mobile servers include ambulances, fire rescue trucks, on-demand delivery systems, and disaster response units. In his seminal work, Larson (1974) introduced a general model for mobile servers, known as the hypercube queueing models. The model incorporated geographical components, or atoms, each with unique arrival rates, and mobile servers with distinct service rates. This model can capture the light traffic regime of police operations, compute the steady-state distribution of binary states, and conveniently evaluate general performance metrics.
The motivation for this study stems from various factors, including modeling police operations with heavy workload (often linked to staff shortages), severe staff shortages, and the observed boundary effect of crime incidents. These observations are supported by data, as illustrated in Figures 1, 1, and 1. Each factor will be discussed in greater detail below.
[…]
The increasing average police response time in Atlanta by month, from 2013 onwards.
[…]
Operations statistics calculated using 911 data from the Atlanta Police Department, from 2013 to 2022, averaged for each of the 80 beats. These figures show that calls are coming in faster than they can be handled by the police, as indicated by the workload in minutes and the hourly rate of the calls.
[…]
The border effect. Burglary incidents reported by the Atlanta police department in 2017 (represented by red dots) clustered near the border between beats (indicated by black lines).
First, modern police operations, often strained by high workload, necessitate a more flexible queueing model than what is provided by the traditional hypercube queueing model. The hypercube model, while elegant in its simplicity, is limited to representing each server’s status as simply “busy” or “idle” and operates under the assumption of a singular queue for the entire system. However, in today’s context, police operations systems frequently face overwhelming demands, leading to a scenario where all servers are constantly busy, which may not be accurately captured by the hypercube model. Therefore, to effectively understand systems with heavier loads, where there could be multiple queues and each queue might exceed a length of one, our objective is to model the distribution of queue lengths for each server. This approach goes beyond the basic “busy” and “idle” states (represented by binary vectors) used in Larson’s original hypercube queueing models, offering a more detailed and realistic representation of contemporary police operations.
Second, the traditional “one-officer-per-beat” operational mode has become increasingly difficult to implement due to severe staff shortages, which currently pose a significant challenge for many police departments, including those in the Atlanta metro area (Cummings 2022). For example, as of January 2023, suburban Atlanta’s Gwinnett County (GA) reportedly “employed 690 officers out of an authorized strength of 939, leaving more than a quarter of sworn officer positions vacant” (Writer 2023). This leads to busier police operations and challenges in assigning at least one officer to each beat each day. To address officer shortages in practice, some municipalities have implemented flexible service regions within their districts, enabling officers to respond to incidents more effectively. This empirical strategy has proven successful. Moreover, by utilizing restricted service regions with some overlap, departments can strike a balance between full flexibility and complete isolation, ensuring that officers can assist neighboring areas when needed without compromising the efficiency of their primary service areas.
Third, implementing flexible service regions with overlapping beats can help address the so-called “border effect.” For example, in Atlanta, we have observed a notable boundary effect in the distribution of police incidents, with crime rates tending to be higher at the peripheries of police beats (refer to Figure 1 for a burglary case example). Additionally, response times at beat boundaries are often longer due to increased travel distances. Police officers also report that more criminal activity occurs near the borders of two beats, resulting in a higher volume of 911 calls in these areas. The boundary effect may contribute to increased crime rates near the edges of designated police zones, underscoring the need for a more efficient allocation of law enforcement resources. This observation suggests that the proposed overlapping beat structure may offer the demonstrated theoretical advantages and potential additional benefits in combating the border effect.
Motivated by these operational realities, this paper introduces a new queueing framework that generalizes the classical hypercube model to accommodate overlapping service regions, individual queue lengths, and flexible dispatch policies. While overlapping patrol strategies have gained traction in practice, there is currently no mathematical framework that systematically evaluates their performance. Our proposed model addresses this gap and offers the following core contributions:
  1. 1.
    Hypercube queueing models for spatial service systems: We introduce a state-space representation that captures per-server queue lengths and structured service overlaps in spatial service systems, allowing analysis of system dynamics under light and heavy traffic.
  2. 2.
    Efficient steady-state computation for large-scale systems: We introduce a hyperlattice structure that, despite its infinite state space, enables tractable and efficient computation through structured sparsity and state truncation techniques.
  3. 3.
    Application to real-world patrol systems: We validate the model using data from Atlanta Police Department operations, demonstrating how overlapping patrol regions can alleviate operational strain and reduce the adverse effects of boundaries. Our framework offers actionable insights into how resource allocation and dispatch policies influence system performance.
[…]
Illustration of the differences in state spaces. Our model represents the status of each server based on the number of servers waiting in its corresponding queue. In contrast, the hypercube model reduces this to a single tail state when all servers are busy. As a result, our model’s state space forms a “hyperlattice”, expanding the original finite-state hypercube structure into an infinite space. While similar hyperlattice structures have been proposed in prior works, our approach utilizes them to represent a different state space.
To achieve our goal, we consider a Markov model with a potentially large state space represented by integer-valued vectors, unlike the hypercube queueing model, which uses binary-valued state vectors as shown in Figure 1. In this model, each server has its own queue, and the integer in the state represents the number of calls waiting in that queue. Crucially, the policy that manages calls in each queue is determined by a hypergraph that reflects the overlapping nature of these regions. The design of the state space results in a sparse transition matrix that allows for transitions between states whose vectors differ by 1 in the $\ell_{1}$ distance (in contrast, the hypercube queueing model considers Hamming distance one transition for binary-valued state vectors). Based on this, we develop an efficient state truncation technique to relate the generalized hypercube queueing model to the known properties of a general birth-and-death process with state-dependent rates. This reduces the detailed balancing equations to a much smaller linear system of equations, enabling efficient computation for steady-state distributions, which are key in evaluating general performance metrics as recognized in Larson’s original work (Larson 1974). We validate the proposed framework by testing it on both synthetic systems, with and without overlapping regions, and applying it to a real police system in collaboration with the Atlanta Police Department (APD). With the adoption of the “umbrella car” strategy by the APD, where multiple officers collaborate to cover more areas due to understaffing and rising crime rates, our study aims to analyze this operational method and help the police evaluate its effectiveness using our model. The results indicate that deploying more police units does not necessarily enhance policing effectiveness. Instead, overlapping police service regions provide a more flexible and dynamic response to varying call frequencies, significantly improving resource allocation efficiency compared to a rigid, non-overlapping structure.
This paper is organized as follows: We begin with a review of relevant literature on spatial queueing systems and Markov models with similar structures. In Section 3, we introduce a novel queueing framework for evaluating districting plans with overlapping regions. Section 3.1 discusses the problem setup and key assumptions. Section 3.2 presents a generalized hypercube queueing model – the hyperlattice model – for capturing system dynamics, while Section 3.3 introduces an efficient method for estimating its steady-state distribution. Section 3.4 introduces two key performance measures using the hyperlattice model. Lastly, we demonstrate the effectiveness of our framework through experiments on synthetic systems in Section 4 and conclude in Section 5, discussing potential extensions and generalizations.
2 Related works
Spatial queueing systems modeling with overlapping service regions has received limited attention in the literature. However, recent developments in service systems and their increasing complexity have sparked a growing interest in this field, particularly in the pursuit of more efficient dispatch policies. For instance, the concept of overlapping zones is discussed in Bammi (1975), which only considers completely overlapping service regions. This approach may need to address higher crime rates near boundaries. In contrast, our study focuses on partially overlapping service regions, which are more challenging to analyze but better suited for our purposes.
Richard Larson’s paper (Larson 1974) and his co-authored book with Asaf Odoni, “Urban Operations Research” (Larson and Odoni 1981), introduced the hypercube queueing model, a pioneering work in this field. The hypercube queueing model provides a spatial queueing framework for analyzing and evaluating the performance of service systems. It permits servers to overlap into each other’s territories and provide assistance when busy, making it particularly useful for emergency service systems, such as police and ambulance operations (Ansari et al. 2017, de Souza et al. 2015, Takeda et al. 2007, Zhu et al. 2020, 2022, Morabito et al. 2008). However, it assumes identical service from all servers sharing a single queue, which may be oversimplified for real-world service systems and is unable to capture more complex systems. In our context, servers may have limited authorized movement, resulting in separate queues for each server. This is because other servers cannot travel to specific regions to offer assistance. For instance, Beojone et al. (2021) proposed an extension of the hypercube model with fully dedicated servers. While their model allows both dedicated and basic servers to serve the same region under certain random dispatch strategies, it differs significantly from our hyperlattice queueing model in how server responsibilities are handled. In their model, dedicated servers are specifically assigned to handle the most severe calls from any region. In contrast, our hyperlattice queueing model assigns equal responsibilities to all servers, focusing on service regions that can be covered by multiple servers. A similar distinction is found in de Souza et al. (2015), where the model considers queue priorities and assigns different responsibilities to servers, such as Advanced Life Support (ALS) units and Basic Life Support (BLS) units. In their models, overlapping service occurs only across different server types. Moreover, Takeda et al. (2007) considered decentralized but dedicated service areas for ambulances, with each ambulance assigned to a specific station and region. In their hypercube model, overlapping service is achieved by enlarging the service regions with decentralized units. In contrast, our hyperlattice model allows units to serve neighboring overlapping regions without requiring decentralization.
Among existing works extending the classic hypercube queueing model for specific applications, a notable contribution is Beojone et al. (2021), which introduces a variation of the hypercube model. Their extension is designed to represent a dispatch policy in which advanced-equipped servers serve solely life-threatening calls (called dedicated servers). Our model, however, differs as it allows servers to operate within the territories of other servers, presenting a distinct problem framework. Additionally, they grouped servers with similar characteristics in their state space representation, aiming to reduce computational demands. In our method, each server is still represented individually in the state space, managing to do so without significant computational expense. We also assume that each server has its own queue, diverging from most hypercube model variations, which typically use a single queue for all servers. A detailed comparison between various hypercube models and our hyperlattice model is presented in Table 1.
Table 1: Comparison of Hyperlattice Model with Hypercube Models
Model Feature Larson (1974) Beojone et al. (2021) Our Model
Service Region Fully overlapping Fully overlapping Flexible overlapping
Call Priority Uniform Tiered (Urgent/Basic) Uniform
Dispatch Policy Priority-based Priority-based Flexible (state-dependent)
Queueing Structure Single loss queue Single general queue Multiple general queue
Tractability High Moderate High (via truncation)
Recently, Tsitsiklis and Xu (2017) introduced a flexible queue architecture for a multi-server model, comprising $n$ flexible servers and $n$ queues connected through a bipartite graph, with each server having its own associated queue. The paper primarily focuses on the model’s theoretical aspects and highlights the key benefits of flexibility and resource pooling in such queueing networks. They specifically concentrate on the scaling regime where the system size $n$ tends to be infinite while the overall traffic intensity remains constant. However, their model cannot be directly extended to our scenario, as it does not account for the geographical relationships among servers.
The queueing literature extensively examines load-balancing models involving multiple dispatchers and flexibility constraints, often under the framework of flexible queues or load-balancing systems. A representative example is the “Join the Idle Queue” policy and related dispatching strategies designed to optimize system performance under varying workload conditions (see, e.g., Cruise et al. (2020), Weng et al. (2020)). These models typically focus on analyzing optimal or near-optimal dispatching policies under simplified assumptions. However, our contribution is primarily modeling-oriented: We develop a generalized hypercube queueing framework that captures the spatial and structural complexity of real-world service systems. Unlike much of the load-balancing literature, which often assumes homogeneous servers, centralized coordination, or simplified system dynamics, our model explicitly incorporates server-level queues, spatially dependent service areas, and flexible, state-dependent dispatch policies. Instead of focusing on specific policies, we develop a general and extensible framework that approximates system performance across a broad class of state-dependent dispatch policies, while maintaining computational tractability.
3 Proposed framework
This section develops a framework for assessing districting plans with overlapping patrol units for service systems. To gain insight into the strategy of overlapping patrol, we consider a moving-server queueing system with multiple servers. Typically, each server (e.g., a patrol service unit) is assigned to one particular service region, where the server has primary responsibility. When a server is not busy serving any active calls, it traverses its home region to perform preventive patrol. A key feature of our model is its consideration of a more flexible state-dependent policy, allowing servers to collaborate and jointly serve a third overlapping service region within a spatial queueing framework, as illustrated in Figure 3.1. For the reader’s convenience, Table EC1 contains summary definitions of notations frequently used in the paper.
3.1 Problem setup and key assumptions
In this section, we describe our problem setup and several key assumptions. Consider a service system patrolled by multiple servers, where each server is allocated to a specific area and has its own queue. Unlike conventional service systems with separate service territories, our approach permits servers to have overlapping service areas. The entire service system can be represented by an undirected hypergraph (Bretto 2013) denoted by $(\mathcal{I},\mathcal{E})$. In this hypergraph, $\mathcal{I}=\{i=1,\dots,I\}$ denotes the set of vertices, each representing an individual server, with $I$ being the total number of servers in the system. The set $\mathcal{E}$ represents the hyperedges, where each hyperedge $e\in\mathcal{E}$ is a subset of $\mathcal{I}$, i.e., $\mathcal{E}\subseteq 2^{\mathcal{I}}$. Each hyperedge corresponds to a group of servers jointly overseeing an overlapping area. The cardinality of $\mathcal{E}$ is denoted by $E$. Figure 3.1 presents the hypergraph of a three-server system as an example.
The service system encompasses a geographical space denoted by $\mathcal{S}\subset\mathbb{R}^{2}$. We denote the part of the space that server $i$ patrols as $\mathcal{S}_{i}\subseteq\mathcal{S}$. We require that the service territories of all servers cover the entire space of the system, i.e., $\bigcup_{i\in\mathcal{I}}\mathcal{S}_{i}=\mathcal{S}$. Regions jointly patrolled by a set of servers $e\in\mathcal{E}$, are termed as overlapping service regions, and are denoted as $\mathcal{O}_{e}$. Conversely, regions exclusively patrolled by server $i$ are termed primary service regions, and are defined as $\mathcal{P}_{i}=\mathcal{S}_{i}\backslash\bigcup_{\{e:i\in e\}}\mathcal{O}_{e}$. Note that it is important to guarantee $e\not\subseteq e^{\prime}$ for every $e,e^{\prime}\in\mathcal{E}$ and $e\neq e^{\prime}$. This ensures all primary and overlapping service regions are mutually exclusive and can be distinctly identified either by $i\in\mathcal{I}$ or by $e\in\mathcal{E}$. Both primary and overlapping service regions can be arbitrarily small to avoid quantization errors, and they can have any geometric shape.
[…]
Examples of 2- and 3-server systems with overlapping patrols. The white region $\mathcal{P}_{i}$ denotes the primary service area of server $i$, which can only be served by that server. The shaded region $\mathcal{O}_{e}$ represents the overlapping service area $e=\{i,j\}$, which can be jointly served by servers $i$ and $j$.
[…]
The hypergraph of a three-server system. The black dots represent vertices and the shaded areas represent hyperedges.
Assume service calls arrive in the system according to a Poisson process with constant arrival rate. The arrivals of calls in primary service region $i$ and overlapping service region $e$ follow time-homogeneous Poisson processes with rate $\lambda_{i}$ and $\lambda_{e}$, respectively. Let $\lambda=\sum_{i\in\mathcal{I}}\lambda_{i}+\sum_{e\in\mathcal{E}}\lambda_{e}$ be the total call arrival rate in the entire region.
The service process is described as follows. The server begins processing the call as soon as it arrives at the location where the call is reported. The service time of each server is independent of the customer’s location and the system’s history. The service time of server $i$ for any call for service has a negative exponential distribution with mean $1/\mu_{i}$. Let $\mu=\sum_{i\in\mathcal{I}}\mu_{i}$ denote the total service rate in the entire region. After completing the service, the servers return to or remain within their primary service territory, regardless of where the previous call originated. For ergodicity, it is necessary that $\lambda<\mu$ should hold (a detailed stability condition is provided in Section 3.2.5).
3.2 Hyperlattice queueing model for overlapping patrol
We propose a generalized hypercube queueing model, called the hyperlattice model, to capture the operational dynamics of systems with overlapping service regions. The system state depends on the status of all the servers in this system, and the number of calls in each queue to be processed.
3.2.1 State representation
These states can be represented by a hyperlattice in dimension $I$. Each node of the hyperlattice corresponds to a state $B=(n_{i})_{i\in\mathcal{I}}$ represented by a tuple of numbers, where a non-negative integer $n_{i}\in\mathbb{Z}_{+}$ indicates the status of server $i$. Server $i$ is idle if $n_{i}=0$ and busy if $n_{i}>0$. The value of $n_{i}-1$ represents the number of calls waiting in the queue of server $i$ when the server is busy. Figure 3.2.1 (a) gives an example of the state space of a hyperlattice queueing model for a system with two servers. It is worth noting that the state space in a hyperlattice queueing model can be divided into two parts, with one part consisting of states that possess at least one available server (referred to as non-saturated states) and the other part comprising states in which all servers are busy (referred to as saturated states), as shown in Figure 3.2.1.
[…]
The state space in a hyperlattice queueing model for a system with two servers. The state space is shown in (a) as an overview, with each state denoted by the pair $(n_{1},n_{2})$. Green arrows indicate downward transitions and red arrows indicate upward transitions. (b) presents the birth-death model by truncating states at an upper bound $K=2$. In (c), (d), and (e), the details of the state space are depicted in three distinct scenarios.
[…]
An illustration of two types of states in hyperlattice queueing models with different numbers of servers. The area colored red signifies the saturated states that have no idle servers, whereas the empty area represents the non-saturated states that have at least one server that is unoccupied.
3.2.2 Touring algorithm
To create the transition rate matrix, the states must be arranged in order. However, since the standard vector space lacks an inherent order, we employ a touring algorithm to traverse the entire hyperlattice. This algorithm generates a sequence of $I$-digit non-negative integers $B_{0},B_{1},\dots$, which contains an infinite number of members and fully describes the order of the states.
[…]
A hyperlattice queueing model for a system with three service regions. (a) shows part of the state space of the hyperlattice for $k=1,2,3$; (b) presents the tour of all the states with $k\leq 3$, where solid arrows indicate transitions within states with the same number of units, while dashed arrows represent transitions between states with different numbers of units; (c) shows the sequence of all states within $k=2$, where the number of balls separated by dashed circles represents the status of server $i$ ($n_{i}$).
Input: Number of servers II; Output: The sequence of ordered states $\mathcal{B}$;
Initialization: $\mathcal{B}\leftarrow\emptyset$;
for $k=0,1,\dots$ do
 $\mathcal{B}_{k}\leftarrow\{(n_{i})\in\mathbb{Z}_{+}^{I}\mid\sum_{i\in\mathcal{I}}n_{i}=k\}$, where $|\mathcal{B}_{k}|=\binom{k+I-1}{I-1}$;
   Sort $\mathcal{B}_{k}$ using an arbitrary order;
 $\mathcal{B}\leftarrow(\mathcal{B},\mathcal{B}_{k})$;
end for
Algorithm 1 Tour algorithm for a hyperlattice
We simplify the tour problem by decomposing it into $K$ sub-problems. Each sub-problem $k\leq K$ first identifies the states in which exactly $k$ calls are staying in the system (either waiting in the queue or being served), and then proceeds to solve the enumerative combinatorics problem by searching for $\binom{k+I-1}{I-1}$ possible combinations in a pre-determined sequence. It is important to note that the sequence generated by each sub-problem is just one of numerous possible tours of the cutting plane in the hyperlattice. Figure 3.2.2 presents a hyperlattice with $I=3$, along with its corresponding tour of all the states. The combinatorics can be translated into the “stars and bars” (Flajolet and Sedgewick 2009), i.e., finding all possible positions of I1I-1 dashed circles that separate $k$ balls into $I$ groups, as illustrated in Figure 3.2.2 (c). The number of balls being separated in group $i$ can be regarded as $n_{i}$. Algorithm 1 summarizes the tour algorithm that generates the sequence $B_{0},B_{1},\dots$. For convenience, we index the states based on the order produced by the tour algorithm. The state index is represented by $u\in\mathcal{U}=\{0,1,2,\dots\}$.
3.2.3 Dispatch policy
We define a class of dispatch policies within a general state-dependent framework. Let $\mathcal{H}$ denote the set of all Markovian dispatch policies. Our state-dependent dispatch policy, $\eta\in\mathcal{H}$, is defined as follows: When a call arrives at location $s\in\mathcal{S}$, we first determine whether $s$ belongs to a primary service region or an overlapping service region. If $s\in\mathcal{P}_{i}$ for some server $i\in\mathcal{I}$, the call is routed directly to server $i$. If $s$ belongs to an overlapping region, i.e., $s\in\mathcal{O}_{e}$ for some server set $e\in\mathcal{E}$, then the call can be served by any server in $e$. In this case, the dispatch policy is characterized by a discrete distribution $\eta_{e,u}=(\eta_{e,u}(i))_{i\in e}$, where $\eta_{e,u}(i)$ denotes the probability of assigning the call to server $i\in e$. The corresponding dispatch policy is denoted as $\eta=\left\{\eta_{e,u}\mid e\in\mathcal{E},u\in\mathcal{U}\right\}$. This broad class of state-dependent policies includes canonical examples such as “Join-the-Shortest-Idle-Queue” in Weng et al. (2020).
[…]
Queue networks with two servers and two queues. (a) both servers are idle, (b) one of the servers is busy, and (c) both servers are busy, where $p$ is the probability of assigning the call to server 1.
A simple example of such a dispatch policy for a two-server system with a single overlapping service region is illustrated in Figure 3.2.3, where the distribution $\eta_{e,u}$ is designed to prioritize the idle server. Specifically, in state $u$, if only one server is idle, the dispatch policy assigns the call exclusively to that server, i.e., $\eta_{e,u}(i)=1$ for the idle server $i$. Otherwise, if both servers are either available or busy, the call is assigned randomly, with $\eta_{e,u}(1)=p$ and $\eta_{e,u}(2)=1-p$ for some fixed probability $p\in[0,1]$.
3.2.4 Transition rate matrix
Now we define the transition rate matrix for the hyperlattice queue $Q=(q_{uv})_{u,v\in\mathcal{U}}$, where $q_{uv}$ ($u\neq v$) denotes the transition rate departing from state $B_{u}$ to state $B_{v}$. Let $u$ represent the index of state $(n_{1},n_{2},\dots,n_{i},\dots,n_{I})$ and let $d_{uv}=\left\lVert B_{u}-B_{v}\right\rVert_{1}$ be the $\ell_{1}$ distance (or Manhattan distance) between two vertices $B_{u}$ and $B_{v}$ on the hyperlattice. We define $u^{i+}$ as the index of state $(n_{1},n_{2},\dots,n_{i}+1,\dots,n_{I})$; note that thus we have $d_{uu^{i+}}=1$ and $B_{u^{i+}}\succ B_{u}$ ($B_{u}$ is elementwise greater than $B_{v}$). Diagonal entries $q_{uu}$ are defined such that $q_{uu}=-\sum_{v:v\neq u}q_{uv}$ and therefore the rows of the matrix sum to zero. There are two classes of transitions on the hyperlattice: upward transitions that change a server’s status from idle to busy or add a new call to its queue; and downward transitions that do the reverse. The downward transition rate from state $B_{u^{i+}}$ to its adjacent state $B_{u}$ is always $q_{u^{i+}u}=\mu_{i}$. The upward transition rate, however, will depend on how regions overlap with each other and the dispatch rule when a call is received, which can be formally defined by the following proposition.
Proposition 3.1 (Upward transition rate)
For an arbitrary state index $u$, the upward transition rate from $B_{u}$ to $B_{u^{i+}}$ is:
$q_{uu^{i+}}=\lambda_{i}+\sum_{e\in\mathcal{E}}\eta_{e,u}(i)\lambda_{e},$ (1)
where $\eta_{e,u}(\cdot)$ is a probability mass function that gives the probability of the new call in the overlapping service region $\mathcal{O}_{e}$ being received by server $i\in e$ at state $B_{u}$. The probability mass function $\eta_{e,u}$ has the following properties:
$\eta_{e,u}(i)$ $=0,$ $~\forall i\notin e,e\in\mathcal{E},u\in\mathcal{U},$ (2)
$\sum_{i\in\mathcal{I}}\eta_{e,u}(i)$ $=1,$ $~\forall e\in\mathcal{E},u\in\mathcal{U}.$ (3)
The following argument justifies the proposition. The first term of (1) suggests that, for an arbitrary server $i$, it must respond to a call received in its primary service region $i$. The second term of (1) means that the server $i$ needs to respond to a call received in the overlapping service region $e$ according to the dispatch policy $\eta_{e,u}$. It is essential to note that the dispatch policy $\eta_{e,u}$ is state-dependent and can be defined based on the user’s specific needs, providing a general framework that is suitable for elucidating both random and deterministic dispatch strategies. For example, Figure 3.2.3 represents a specific case of the dispatch policy, in which the assignment of server $i$ to the call received in overlapping service region $e$ can be discussed in two scenarios: (a) When all servers $i\in e$ are busy ($u:n_{i}>0,\forall~n_{i}\in B_{u}$), this call will join the queue of server $i$ based on the probability $\eta_{e,u}(i)$. (b) Conversely, when there is at least one server idle ($u:n_{i}=0,n_{i}\in B_{u}$), the server $i$ will be assigned to this call based on the probability $\eta_{e,u}(i)$.
3.2.5 Markov chain representation
The hyperlattice model introduced above can be formulated as a continuous-time Markov chain (CTMC) on a discrete state space. Specifically, it is characterized by the state space $B$ and the transition rate matrix $Q$. The original state space is defined as $B:=\mathbb{Z}_{+}^{|\mathcal{I}|}$. For convenience, we introduce a touring algorithm that establishes a one-to-one mapping from the high-dimensional state space $B$ to a one-dimensional index state space $\mathcal{U}:=\mathbb{Z}_{+}$. Additionally, let $e_{i}$ denote a vector with all elements zero except for the $i$-th entry, which is set to 1. The transition rate matrix $Q=(q_{uv})_{u,v\in\mathcal{U}}$ is then given by:
$q_{uv}=\begin{cases}\lambda_{i}+\sum_{e\in\mathcal{E}}\eta_{e,u}(i)\lambda_{e}&\text{if }B_{v}=B_{u}+e_{i},\quad\forall i\in I,\\ \mu_{i}&\text{if }B_{v}=B_{u}-e_{i},\left(B_{u}\right)_{i}>0,\quad\forall i\in I,\\ -\lambda-\sum_{i\in I:\left(B_{u}\right)_{i}>0}\mu_{i}&\text{if }B_{v}=B_{u},\\ 0&\text{otherwise.}\end{cases}$ (4)
For the hyperlattice queueing model to be stable, i.e., to attain a steady-state distribution, the corresponding CTMC must be ergodic. To ensure ergodicity, we impose the following stability condition for general policies.
{assumption}
[…]
$\sum_{e\in\mathcal{A}}\lambda_{e}<\sum_{i\in\mathcal{I}:\,\exists\,e\in\mathcal{A},(i,e)\in\tilde{\mathcal{E}}}\mu_{i},\quad\forall\,\mathcal{A}\subseteq\mathcal{N}.$
Assumption 3.2.5 is analogous to Hall’s marriage theorem (Hall 1987). When this condition holds, by max-flow/min-cut theorem (Ford Jr and Fulkerson 1956), there exists flows $\Lambda_{ie}\in\mathbb{R}_{+}^{|\mathcal{I}|\times|\mathcal{N}|}$ with
$\sum_{i\in\mathcal{I}}\Lambda_{ie}=\lambda_{e}\quad(\forall e\in\mathcal{N}),\quad\sum_{e\in\mathcal{N}}\Lambda_{ie}\leq\mu_{i}-\delta_{i}\quad(\forall i\in\mathcal{I}),$
for some $\delta_{i}>0$ and $\Lambda_{ie}=0$ whenever $(i,e)\notin\tilde{\mathcal{E}}$. Then, we define a state-independent policy $\eta$, under which each arrival from primary or overlapping region $e\in\mathcal{N}$ is assigned to server $i\in\mathcal{I}$ according to a fixed probability distribution:
$\eta_{e}(i)=\frac{\Lambda_{ie}}{\sum_{i^{\prime}\in\mathcal{I}}\Lambda_{i^{\prime}e}},\quad\forall i\in\mathcal{I},e\in\mathcal{N}.$ (5)
Lemma 3.2
Under Assumption 3.2.5, there exists a state-independent policy $\eta$ as defined in (5) that ensures stability of the queueing system.
The proof of Lemma 3.2 follows directly from Assumption 3.2.5 and the max-flow/min-cut theorem. Note that this stability condition applies generally to state-independent policies. More sophisticated state-dependent policies (e.g., Join-the-Idle-Queue) can also be designed to ensure stability whenever Hall’s condition is satisfied.
3.2.6 Balance equations
To obtain the steady-state probabilities of the system, we can solve the balance equations given the transition rate matrix $Q$. Let $\mathbb{P}\{B\}$ denote the probability that the system is occupying state $B$ under steady-state conditions. The balance equations can be written as:
$\left(\sum_{v:{B_{v}\succ B_{u}\atop d_{uv}=1}}q_{uv}+\sum_{v:{B_{u}\succ B_{v}\atop d_{uv}=1}}q_{uv}\right)\mathbb{P}\{B_{u}\}=\sum_{v:{B_{u}\succ B_{v}\atop d_{vu}=1}}q_{vu}\mathbb{P}\{B_{v}\}+\sum_{v:{B_{v}\succ B_{u}\atop d_{vu}=1}}q_{vu}\mathbb{P}\{B_{v}\},\quad u\in\mathcal{U}.$ (6)
We also require that the probabilities sum to one, namely,
$\sum_{u\in\mathcal{U}}\mathbb{P}\{B_{u}\}=1.$
However, one can observe that the above balance equations are hard to solve as the number of states on the hyperlattice is infinite and grows exponentially with the number of servers, even when the queue capacity is limited. Additionally, from Equation (4), the transition rates $q_{u,v}$ explicitly depend on the state-dependent dispatch policy $\eta_{e,u}$ for all $e\in\mathcal{E}$ and $u,v\in\mathcal{U}$. This state-dependent structure further complicates solving Equations (6), rendering them computationally intractable. To overcome this difficulty, we introduce an efficient approximation through a truncated hyperlattice model in the following section.
3.3 A truncated hyperlattice model
We approximate the solution to Equations (6) by solving the balance equations for a truncated hyperlattice model, in which the total capacity across all servers is bounded by an upper bound $K$. Specifically, instead of modeling each server as an infinite-capacity queue, we model the entire system analogously to an aggregated Erlang loss queue, i.e., the arrival rate in any region is set to zero whenever the total number of units in the system reaches $K$. In a light-traffic regime, setting the upper bound $K$ sufficiently large allows the truncated hyperlattice model to approximate the stationary dynamics of the original infinite-capacity queues closely. Figure 3.2.1(b) illustrates an example of a 2-server overlapping queueing system truncated at $K=2$, with the original system shown in Figure 3.2.1(a).
3.3.1 Approximate stationary distribution under truncation
Rather than attempting to calculate the stationary distribution across an infinite number of states in the original state space, the problem can be simplified by focusing only on a finite subset of states. These states are selectively considered if their $k$ values are less than a user-defined hyper-parameter $K$, or $\left\lVert B\right\rVert_{1}\leq K$, and the sum of their steady-state probabilities remains to be one. It is worth noting that the stationary probability of the “tail” of the hyperlattice (states with large $k$ value) is usually low and can be negligible since their chance of happening decreases with the length of the queue. Specifically, in a light-traffic regime (e.g., when the arrival rate is significantly smaller than the service rate), the steady-state probability in the tail of the hyperlattice decreases fast as $K$ increases. As a result, the truncated queueing system provides a close approximation of the original system when $K$ is sufficiently large, where the stationary tail probability is negligible in the original system. In contrast, in a high-traffic regime, the truncated hyperlattice model may diverge significantly from the original model with infinite capacity. However, for certain performance metrics, e.g., workload variance across servers, the truncated model can still serve as a reasonable approximation of the original model. The accuracy of this approximation in a high-traffic regime depends primarily on how sensitively the metric of interest responds to the system state. In the following experiments, we demonstrate that when focusing on server workload, the truncated model remains a good approximation of the simulation results.
We represent this set of truncated states by $\mathcal{U}_{K}=\{0,\dots,U_{K}\}$, where $U_{K}$ denotes the size of the set. In the subsequent discussion, Lemma 3.3 presents the equation for $U_{K}$ in relation to $K$ and $I$. This demonstrates that the size of the set can expand exponentially relative to the number of servers $I$, and it can become quite substantial even for moderate values of $K$, particularly in systems with multiple servers.
Lemma 3.3
The size of the set $\mathcal{U}_{K}$ is $U_{K}=\binom{K+I}{I}$.
Proof 3.4
Proof of Lemma 3.3 The size of the set $\mathcal{U}_{K}$ is equal to the summation $U_{K}=\sum_{k=0}^{K}\binom{k+I-1}{I-1}$, which can be simplified using the Hockey Stick identity Jones (1996):
$\sum_{k=n}^{m}\binom{k}{n}=\binom{m+1}{n+1}$.
Applying the identity with m=I1m=I-1 and n=K+I1n=K+I-1, we get:
$\sum_{k=0}^{K}\binom{k+I-1}{I-1}=\sum_{k=I-1}^{K+I-1}\binom{k}{I-1}=\binom{K+I}{I}$
Therefore, the result of the sum is $\binom{K+I}{I}$.
Denote the steady-state probabilities and transition rate matrix for the states in set $\mathcal{U}_{K}$ as $\pi_{K}=\{\mathbb{P}\{B_{u}\}\}_{u\in\mathcal{U}_{K}}$ and $Q_{K}=(q_{uv})_{u,v\in\mathcal{U}_{K}}$, respectively. We can express the balance equations in (6) as a compact form:
$Q_{K}^{\top}\pi_{K}=0,$ (7)
where the right-hand side is a column vector of zeros. The sum of probabilities in set $\mathcal{U}_{K}$ is equal to 1, i.e.,
$\sum_{u\in\mathcal{U}_{K}}\mathbb{P}\{B_{u}\}=1.$ (8)
Therefore, we can compute $\pi_{K}$ by solving (7) and (8). However, there is redundancy in these equations as the number of equations is $K+1$ and the number of unknown variables is $K$. To eliminate this redundancy, we replace one row of $Q_{K}$ with a row of ones, usually the last row, and denote the modified matrix as $Q^{\prime}_{K}$. Similarly, we modify the “solution” vector, which was all zeros, to be a column vector with one in the last row and zeros elsewhere, and denote it as $e_{K}$. In practice, we solve the resulting equation:
$Q_{K}^{\prime\top}\pi_{K}=e_{K}.$
[…]
Support of the transition rate matrix $Q_{K}$ of hyperlattice queueing models (without diagonal entries). The shaded entries correspond to the matrices’ non-zero entries, and the white spaces are zero.
We note that the solution to the above balance equations requires a matrix inversion of $Q_{K}$ (or $Q^{\prime}_{K}$), which can become computationally challenging for large matrix sizes. By Lemma 3.3, the size of $Q_{K}$ can reach a staggering $U_{K}\approx(K+I)^{I}/I!$ (using Stirling’s approximation (Romik 2000)) even for moderate values of $K$ when the system has $I>2$ servers, presenting a considerable computational hurdle. Fortunately, the sparsity of transition rate matrices (as illustrated in Figure 3.3.1) can be leveraged to iteratively find the steady-state distribution using the power method (Nesterov and Nemirovski 2015), by constructing a probability transition matrix of a discrete-time Markov chain through uniformization of the original continuous-time Markov chain. Let $\mathbb{I}$ represent the identity matrix and denote the maximum absolute value on the diagonal of the transition rate matrix $Q_{K}$ as $\gamma=\max|q_{uu}|$. By initializing with a vector $\pi_{K}^{(0)}$, we can compute $\pi_{K}^{(t)}=\pi_{K}^{(t-1)}(\mathbb{I}+Q_{K}/\gamma)$ at each iteration $t=1,2,\dots$, until the distance between $\pi_{K}^{(t)}$ and $\pi_{K}^{(t-1)}$ falls below a specified threshold. The convergence of this approach is guaranteed, and the steady-state distribution can be efficiently computed despite the potentially large size of the transition rate matrix.
3.3.2 Analysis of two canonical policies
Although the stationary distribution of a general state-dependent policy is often intractable to derive, two canonical policies admit explicit characterizations: the Static Random-Routing policy and the Join-the-Idle-Queue policy. We analyze these policies under the truncated hyperlattice model and show that the truncated approximation provides a valid method in the light-traffic regime.
Static Random-Routing policy.
Consider a state-independent static random-routing policy $\eta$. Specifically, recall that $\eta_{e}$ is a vector denoting the probability distribution over servers that determines how an arrival from region $e$ is routed. Then, the system is equivalent to $|\mathcal{I}|$ independent $M/M/1$ queues, where the arrival rate to server $i$ is $\lambda_{i}+\sum_{e\in\mathcal{E}}\eta_{e}(i)\lambda_{e}$ and the service rate is $\mu_{i}$. Assume the stability condition holds, i.e., $\lambda_{i}+\sum_{e\in\mathcal{E}}\eta_{e}(i)\lambda_{e}<\mu_{i}-\delta_{i}$ for each $i\in\mathcal{I}$ and some $\delta_{i}>0$. Then, for server $i$, the stationary distribution of the number-in-system is the geometric distribution $\pi_{i}(n)=(1-\rho_{i})\rho_{i}^{n}$, where $\rho_{i}=\frac{\lambda_{i}+\sum_{e\in\mathcal{E}}\eta_{e}(i)\lambda_{e}}{\mu_{i}}$ is the traffic intensity at server $i$. Then, let $N_{i}$ be a random variable with distribution $\pi_{i}$, and define $S:=\sum_{i\in\mathcal{I}}N_{i}$ as the total number-in-system across all servers. The stationary probability of the truncated state space $\mathcal{U}_{K}$ is $\mathbb{P}\!\left\{S\leq K\right\}$. The probability generating function of $N_{i}$ is
$G_{i}(z)=\mathbb{E}\left[z^{N_{i}}\right]=\frac{1-\rho_{i}}{1-\rho_{i}z},\quad 0\leq z<\frac{1}{\rho_{i}}.$
Since each server is independent, the probability generating function of $S$ is $\prod_{i\in\mathcal{I}}G_{i}(z)$. Then, for any $z\in\left(1,1/\rho_{\max}\right)$ with $\rho_{\max}:=\max_{i}\rho_{i}$, Markov’s inequality on $z^{S}$ yields
$\mathbb{P}\{S\geq K\}=\mathbb{P}\left\{z^{S}\geq z^{K}\right\}\leq z^{-K}\prod_{i\in\mathcal{I}}\frac{1-\rho_{i}}{1-\rho_{i}z}\leq C(z)e^{-K\ln z},$
where $C(z):=\prod_{i\in\mathcal{I}}\frac{1-\rho_{i}}{1-\rho_{i}z}$ is a finite constant given $z$. Then, for any $z\in\left(1,1/\rho_{\max}\right)$,
$\sum_{u\in\mathcal{U}_{K}}\mathbb{P}\left\{B_{u}\right\}=\mathbb{P}\{S\leq K\}=1-\mathbb{P}\{S\geq K+1\}\geq 1-C(z)e^{-(K+1)\ln z}.$ (9)
Eq. 9 establishes that the stationary probability of the truncated state space converges to 1 exponentially fast in $K$. As an illustrative example of a light-traffic two-server system with $\rho_{1}=\rho_{2}=1/2$ and $z=3/2$, we find that $C(z)=4$, and hence $\sum_{u\in\mathcal{U}_{K}}\mathbb{P}\{B_{u}\}\geq 1-4e^{-(K+1)\ln(3/2)}=1-4\left(\tfrac{2}{3}\right)^{K+1}.$ In particular, when taking $K=10$, the stationary probability of the truncated state space $\mathcal{U}_{K}$ exceeds $0.95$. This example demonstrates that the truncated hyperlattice model introduced in Section 3.3.1 provides a valid and accurate approximation in the light-traffic regime.
Join-the-Idle-Queue policy.
Consider a Join-the-Idle-Queue policy in which each arrival, if possible, is routed uniformly at random to one of the idle servers. When all servers are either busy or all idle, the policy reduces to a static random-routing policy $\eta$. The dynamics under this policy are more intricate, as they are state-dependent and must distinguish whether each queue is empty or not. However, this system can be coupled with the one operating under the pure static random-routing policy $\eta$. In particular, let $S^{\ast}(t)$ denote the total number-in-system process (i.e., a birth-death process) under the Join-the-Idle-Queue policy, and let $S(t)$ denote the corresponding coupled process under the pure random-routing policy $\eta$. One can show that $S^{\ast}(t)\leq S(t)$ almost surely for any time $t>0$, since both processes have identical arrival (birth) rates, but under the Join-the-Idle-Queue policy the total service (death) rate is never smaller than under static random routing, as it ensures that at least as many servers remain busy at all times. Therefore, $S^{\ast}$ is stochastically dominated by $S$, implying that the stationary probabilities satisfy $\mathbb{P}\{S^{\ast}<K\}\leq\mathbb{P}\{S<K\}$ (see, e.g., classical results on stochastic comparison of birth-death processes in Van Doorn (2012)). That is, the tail probability bound in Eq. (9) also applies to $S^{\ast}$ under the Join-the-Idle-Queue policy, and the truncated hyperlattice model remains a valid approximation under the light-traffic regime.
3.3.3 Generalization of the hypercube model
The hyperlattice queueing model serves as a natural generalization of Larson’s hypercube model, introduced initially in Larson (1974). Specifically, when the entire region is treated as a single overlapping service region for all servers, i.e., $\mathcal{E}=\{\mathcal{I}\}$, the hyperlattice model reduces to the hypercube model under a given dispatch policy. In the hypercube model, the dispatch policy $\eta$ follows a “fixed preferences” rule, where the dispatcher always selects the highest-ranked available unit based on a predefined preference list, such as the closest-distance dispatching policy. Additionally, the basic hypercube model in Larson (1974) is formulated with each server as an Erlang loss queue with zero capacity. This approach is analogous to our truncated hyperlattice model, except that we extend the loss framework to the entire system rather than individual servers.
Beyond these similarities, our hyperlattice queueing model extends the hypercube model in two fundamental ways. First, it differentiates between overlapping and primary service regions, whereas the hypercube model assumes a single, fully overlapping region. Second, it enables a more flexible, state-dependent dispatch policy that accounts for overlapping services. In the experiments in EC2, we demonstrate that the hyperlattice queueing model provides a more accurate approximation than the hypercube model under complex queueing dynamics with overlapping services.
3.4 Measures of performance
In this section, we explicate two key performance measures for a service system with overlapping service regions, using the proposed hyperlattice queueing model as a means of analysis.
3.4.1 Individual workloads
The workload of server $i$, denoted by $\rho_{i}$, is simply equal to the fraction of time that server $i$ is busy serving calls. Thus $\rho_{i}$ is equal to the sum of the steady-state probabilities on part of the hyperlattice corresponding to $n_{i}>0$, i.e.,
$\rho_{i}=\sum_{u\in\mathcal{U}}\mathbbm{1}\{u:n_{i}>0,n_{i}\in B_{u}\}\cdot\mathbb{P}\{B_{u}\},$ (10)
where $\mathbbm{1}$ denotes indicator function and $n_{i}$ represents the status of server $i$ in state $B_{u}$. In practice, because the efficient model estimation derived in Section 3.3 only computes the steady-state probabilities for the states in $\mathcal{U}_{K}$, the workload of server $i$ is numerically approximated by
$\rho_{i}\approx\sum_{u\in\mathcal{U}_{K}}\mathbbm{1}\{u:n_{i}>0,n_{i}\in B_{u}\}\cdot\mathbb{P}\{B_{u}\}.$
The individual workloads can be further used to calculate various types of system-wide workload imbalance defined in Larson (1974).
3.4.2 Fraction of dispatches
For the remainder of the system performance characteristics, it is necessary to compute the fraction of dispatches that send a server to each sub-region under its responsibility. We use $\rho_{i,e}$ to represent the fraction of dispatches from sending server $i$ to one of its primary overlapping service regions $\mathcal{O}_{e}$. We have
$\rho_{i,e}=\begin{cases}\sum_{u\in\mathcal{U}}\eta_{e,u}(i)\lambda_{e}\mathbb{P}\left\{B_{u}\right\}/\lambda,&\text{ (a) }i\in e,\\ 0,&\text{ (b) }i\notin e,\end{cases}$ (11)
In addition, we use $\rho_{i,i}$ to represent the fraction of dispatches that send the server $i$ to its primary service region $i$, which is consistently equal to $\lambda_{i}/\lambda$ since this region can exclusively be handled by server $i$. We note that $\rho_{i,e}=0$ if $i\notin e$ and $\sum_{i,e}\rho_{i,e}+\sum_{i}\rho_{i,i}=1$. Additionally, in practice, we consider $u\in\mathcal{U}_{K}$ rather than $\mathcal{U}$ in (11) as we only calculate the steady-state probabilities for the first $U_{K}$ states according to Section 3.3.
3.4.3 Travel time estimation
The proportion of dispatches can be employed to derive expressions for travel time metrics. These metrics can be obtained similarly according to Larson (1974). For any server $i\in\mathcal{I}$ and one of its primary overlapping service regions with server set $e\in\mathcal{E}$, let $t_{i,e}$ denote the mean travel time for sending server $i$ to the service region $\mathcal{O}_{e}$, and let $t_{i,i}$ denote the mean travel time for sending server $i$ to its primary service region. The unconditional mean travel time, denoted as $\bar{\tau}$, can be estimated by
${\bar{\tau}}=\sum_{i\in\mathcal{I}}\left(\sum_{e\in\mathcal{E}}\rho_{i,e}t_{i,e}+\rho_{i,i}t_{i,i}\right).$ (12)
Similarly, the mean travel time for each individual server $i\in\mathcal{I}$, denoted as $\tau_{i}$, can be estimated by
${\tau}_{i}=\frac{\sum_{e\in\mathcal{E}}\rho_{i,e}t_{i,e}+\rho_{i,i}t_{i,i}}{\sum_{e\in\mathcal{E}}\rho_{i,e}+\rho_{i,i}},$ (13)
and the mean travel time for each overlapping service region $e\in\mathcal{E}$, denoted as $\tau_{e}$, is given by
${\tau}_{e}=\frac{\sum_{i\in\mathcal{I}}\rho_{i,e}t_{i,e}}{\sum_{i\in\mathcal{I}}\rho_{i,e}}.$ (14)
However, it is important to note that this approach assumes that queueing delay is the predominant factor influencing the server’s response time for each incident. This assumption has also been discussed in the hypercube queueing model by Larson (1974). Moreover, in this metric, travel times are not explicitly embedded within the model formulation. Instead, following Larson’s hypercube approach, travel times are treated as exogenous factors and assessed through post-hoc analyses.
4 Experiments
In this section, we evaluate the performance of the hyperlattice model using the proposed metrics in Section 3.4. Our assessment includes a simulated environment in Section 4.1 and a real-world case study on police redistricting in Atlanta, Georgia, conducted in collaboration with the Atlanta Police Department in Section 4.2.
4.1 Synthetic results
We first evaluate the effectiveness of the proposed hyperlattice queueing model through synthetic experiments. These experiments employ diverse parameter settings within synthetic service systems to investigate the accuracy of the performance metrics estimated by our models. To embark on this, we first use the hyperlattice queueing model to estimate the steady-state probabilities of the systems by solving the balance equations (6). Based on the solutions, we then calculate summary statistics, including individual workloads and the fraction of dispatches, as shown in (10) and (11). To validate the accuracy of our model estimations, we also developed a simulation program using simpy (Prountzos et al. 2021), a process-based discrete-event simulation framework based on Python. The primary purpose of this simulation program is to provide a ground truth for the queueing dynamics of the system, allowing us to compare the estimated performance measures from our model with the simulation results.
The simulation setup involves a $100\times 100$ grid, where the width and height are both set to $100$ units. We explore different levels of overlapping service areas by varying the overlap width from 0 to $99$ in $100$ increments. For each overlap width, we calculate the subregion radius and set the positions of the two servers accordingly. The servers are positioned symmetrically along the $x$-axis at coordinates determined by the radius of the subregion, and each server is responsible for a specific subregion polygon. When an event occurs, its location is randomly determined within the region. The simulation then identifies the police units responsible for the event based on their assigned regions. If multiple police units are available, a dispatch strategy is employed to select which unit will respond. The selected police unit travels to the event location, and the travel time is computed based on the Euclidean distance between the unit’s current position and the event location. Upon arrival, the police unit handles the event, which involves a service time drawn from an exponential distribution based on the unit’s service rate. The police units have a service rate of $1$ and move at a constant speed of 10 units, ensuring that queuing delay remains the predominant factor influencing the server’s response time for each incident. The event arrival rate is set to $1$, with each simulation run lasting $1000$ events. Using the simulation results, we can approximate the true individual workload of each server $\rho_{i}$ by calculating the percentage of time that the server is busy, and the true fraction of dispatches $\rho_{i,e}$ (or $\rho_{i,i}$) by calculating the percentage of calls received in region $e$ (or $i$) that have been assigned to server $i$.
The experiments were conducted on a personal laptop running macOS, equipped with an M3 Pro chip featuring a $12$-core CPU with a clock speed of up to $4.06$ GHz and $18$ GB of LPDDR5 RAM with a memory speed of up to $153.6$ GB/s.
4.1.1 Overlapping patrolling
We also test our models on the synthetic systems that allow overlapping patrolling. Adapting the truncated hyperlattice approximation from Section 3.3 with an upper bound state of $K=10$, we create a two-server system where the service region is a 100 by 100 square with a single overlapping service region in the middle ($I=2$ and $E=1$), as depicted in Figure 3.1 (a). We assume that calls are uniformly distributed with an arrival rate of $\lambda=1$ and that the service rate of each server is fixed at $\mu_{i}=1$. Changing the hyperparameters of the system, such as the size of the overlapping region and the dispatch policy, can significantly shift the queueing dynamics. Our numerical experiments suggest that the hyperlattice queuing models can accurately capture intricate queuing dynamics under different hyperparameter settings. In Figures EC2 (a) and (b), we observe that the approximation error of both the individual workloads and the fraction of dispatch by the hyperlattice model is less than $0.1$ across all $\lambda/\mu$ ratios. Additionally, in Figure EC2 (c), we show that the mean response time of the system is also closely approximated by the hyperlattice model.
[…]
Synthetic results for two-server systems with varying overlapping ratio $r$. In (a), (b), and (c), the shaded area represents the overlapping region of the system. (d) and (e) show individual workloads $\rho_{i}$ and fraction of dispatches $\rho_{i,(i,j)}$ with varying overlapping ratio $r$. The solid lines represent the estimated performance measures using hyperlattice queueing models, whereas the dashed lines represent the true performance measures obtained from simulation runs.
To examine the impact of the overlapping ratio on system dynamics, we first create 100 systems by varying the overlapping ratio $r$. This ratio represents the percentage of the area of the overlapping region with respect to the area of the entire service region of server 1, i.e., $r=|\mathcal{O}_{\{1,2\}}|/(|\mathcal{O}_{\{1,2\}}|+|\mathcal{P}_{1}|)$, as shown in Figure 4.1.1 (a-c). We then compare the estimated values of $\rho_{i}$ and $\rho_{i,e}$ to their true values (indicated by dashed lines) in Figure 4.1.1 (d) and (e), and find that the estimated values are a good approximation of their true counterparts. Clearly, the workloads of the two servers differ more as the overlapping ratio increases, thereby enlarging the area that server 1 has to cover. Moreover, as the ratio increases, the probability of dispatching either server to the overlap region grows (indicated by green and orange lines). In contrast, the probability of sending server 2 to its primary area remains the same (blue line), and the chance of dispatching server 1 decreases (red line). This observation arises from our simplified synthetic setup, where we use a uniform dispatch policy and the dispatch probability to a region directly correlates with its size.
[…]
Synthetic results for two-server systems with varying dispatch policy $\eta_{(1,2)}(1)$ and $\eta_{(1,2)}(2)$. In (a), (b), and (c), the shaded area represents the overlapping region of the system. (d) and (e) show individual workloads $\rho_{i}$ and fraction of dispatches $\rho_{i,(i,j)}$ with varying overlapping ratio $r$. The solid lines represent the estimated performance measures using hyperlattice queueing models, whereas the dashed lines represent the true performance measures obtained from simulation runs.
Furthermore, we also investigate the impact of dispatch policy by creating another 100 systems with different values of $\eta_{(1,2)}(1)$ and $\eta_{(1,2)}(2)$ (where $\eta_{(1,2)}(1)=1-\eta_{(1,2)}(2)$), while keeping the service regions the same, as shown in Figure 4.1.1 (a-c). For simplicity, we here omit the subscript state $u$ in the notation of dispatch policy $\eta_{e,u}(i),~\forall i\in e$ as we assume it depends on the state $u$ only only when prioritizing idle servers and remains independent of $u$ when all servers are busy, as discussed in Section 3.2.3. Figure 4.1.1 (d) and (e) show that the estimated $\rho_{i}$ and $\rho_{i,e}$ (indicated by solid lines) can closely approximate their true values suggested by the simulation. As we observe, when the dispatch probability for a call from the overlapping region to either server is 0.5, the workload distribution between the two servers is equal, i.e., $\eta_{1,2}(1)=\eta_{1,2}(2)=0.5$.
4.2 Case study: Police Redistricting in Atlanta, GA
[…]
Police workload in Atlanta, Georgia. (a) presents a comprehensive map delineating the various police zones in Atlanta, Georgia, while (b) specifically focuses on Zone 5. The intensity of color shading directly corresponds to the frequency of police call arrivals, providing a visual representation of call distribution across the zones.
In this section, we present a case study focusing on the police redistricting problem within Atlanta, Georgia. Figure 4.2 illustrates the various police zones of the city, with a particular emphasis on the intricate beats of Zone 5. Ideally, each beat would be served by a dedicated police unit, ensuring a focused response within its boundaries. However, this is often not feasible due to a limited number of available police units, primarily attributed to constraints in officer and resource allocation. In instances where a specific beat lacks a dedicated police unit, units from adjacent beats are required to assist. While such support is permissible among units within the same zone, assistance across zones is explicitly prohibited. The depth of color shading in Figure 4.2 is a direct representation of the frequency of police call arrivals, quantified by the arrival rate $\lambda$. These rates are estimated based on an analysis of historical 911 calls-for-service data, specifically collected during the years 2021 and 2022, as detailed in the study of Zhu et al. (2022). The color indicates that Zone 5 experiences the highest frequency of police calls, averaging approximately 18 calls per hour. As a result, our analysis concentrates on examining potential overlapping districting plans specific to the beats in Zone 5. A key performance metric under consideration is the standard deviation of the individual workloads, as previously discussed in Section 3.4.1.
[…]
An illustrative example showing different districting plans for Zone 5 in Atlanta. Gray lines represent the basic geographical units patrolled by the police, and red lines outline the districting plans. In the map, regions shaded in grey with dashed blue lines represent beats without a dedicated police unit. The presence of blue arrows originating from these beats signifies that their policing needs are addressed by units from adjacent beats.
In Figure 4.2, we measure the standard deviation of the individual workloads for different overlapping districting plans using the truncated hyperlattice model in Section 3.3 with an upper bound state of $K=10$, denoted as $\sigma$. Panels (a), (b), and (c) display a scenario where a single beat lacks a dedicated police unit. Specifically, in panel (a), the districting plan is non-overlapping since the policing requirements of this unassigned beat are exclusively managed by a single adjacent beat. In panels (b) and (c), we display different overlapping districting plans by varying the number of overlapping beats. In these configurations, the standard deviations of workload are notably reduced, measured at $0.11$ and $0.08$ respectively, demonstrating the impact of overlapping districts on workload distribution uniformity. Panels (d), (e), and (f) depict scenarios wherein two beats operate without dedicated police units. Notably, panel (f) demonstrates a further reduction in workload standard deviation compared to panel (c), measured at $0.06$, attributable to an additional beat being serviced by the overlapping police units from its neighbors.
Table 2: Workload allocation among police units across beats under various districting plans.
Plan Workload of police units $\sigma$
1 2 3 4 5 6 7 8 9 10 11 12
a 0.28 0.63 - 0.19 0.27 0.19 0.19 0.26 0.27 0.23 0.13 0.21 0.13
b 0.51 0.46 - 0.21 0.28 0.21 0.20 0.28 0.28 0.25 0.14 0.22 0.11
c 0.41 0.35 - 0.33 0.40 0.21 0.21 0.28 0.29 0.25 0.14 0.22 0.08
d 0.27 0.63 - 0.19 0.26 0.19 0.19 0.26 0.26 0.36 - 0.20 0.13
e 0.51 0.46 - 0.20 0.28 0.21 0.20 0.34 0.28 0.32 - 0.22 0.10
f 0.41 0.35 - 0.33 0.40 0.21 0.24 0.31 0.29 0.29 - 0.26 0.06
g 0.29 0.23 0.45 0.21 0.28 0.21 0.20 0.28 0.28 0.25 0.13 0.22 0.07
In Table 2, we record the workload allocation among police units across beats under different districting plans. Plans (a) – (f) align with corresponding panels in Figure 4.2. Plan (g) depicts an independent scenario where each beat is serviced by a dedicated police unit with non-overlapping districts. It is noteworthy that the total number of police units varies by districting plan, with plans (a) – (c) utilizing 11 units, plans (d) – (f) utilizing 10 units, and plan (g) utilizing all 12 units. We observe that plan (f) exhibits a lower standard deviation in workload compared to plan (g), despite the latter having two additional police units.
Table 3: Mean travel times for police units across beats under different districting plans.
Plan Mean travel time of police units (minutes) $\bar{\tau}$}
1 2 3 4 5 6 7 8 9 10 11 12
a 12.15 14.24 - 12.26 12.07 12.20 11.81 11.94 12.01 12.09 12.16 11.82 12.54
b 13.89 13.64 - 12.26 12.07 12.20 11.81 11.94 12.01 12.09 12.16 11.82 12.59
c 13.24 13.10 - 13.75 12.62 12.20 11.81 11.94 12.01 12.09 12.16 11.82 12.62
d 12.15 14.24 - 12.26 12.07 12.20 11.81 11.94 12.01 12.69 - 11.82 12.61
e 13.90 13.64 - 12.26 12.07 12.20 11.81 12.41 12.01 12.44 - 11.82 12.67
f 13.24 13.10 - 13.75 12.62 12.20 12.21 12.19 12.01 12.28 - 11.99 12.63
g 12.15 12.03 12.21 12.26 12.07 12.20 11.81 11.94 12.01 12.09 12.16 11.82 12.07
In Table 3, we present the mean travel times of different police units under various districting plans, as well as the overall mean travel time $\bar{\tau}$}. It is observed that enabling overlapping service regions results in increased mean travel times for police units serving neighboring areas. Although plan (f) exhibits a lower standard deviation in workload compared to plan (g) with fewer police units, it increases the mean travel time for neighboring units by more than one minute and the overall mean travel time by about half a minute.
The case study also provides valuable managerial insights for addressing the police redistricting problem. It shows that deploying a larger number of police units does not necessarily equate to more effective policing. For instance, plan (f) with 10 units had a lower standard deviation of workload compared to plan (g) with 12 units. This suggests that utilizing overlapping police beats, as opposed to a rigid, non-overlapping structure, can significantly enhance the efficiency of resource allocation. However, it should be noted that utilizing overlapping police beats may come at the cost of increased mean travel times. Practitioners must carefully balance the trade-off between response time and resource allocation efficiency in their actual implementations. In general, the overlapping districting approach is preferred in high-demand areas, such as Zone 5 in Atlanta or other cities facing similar challenges.
5 Discussions
This paper presents a novel generalized hypercube queueing model that captures the detailed queueing dynamics of each server in a system. We demonstrate the model’s versatility by applying it to analyze overlapping patrols. Our Markov model employs a hyperlattice to represent an extensive state space, where each vertex corresponds to an integer-valued vector. The transition rate matrix is highly sparse, permitting transitions between vectors with an $\ell_{1}$ distance difference of 1. To enhance computational tractability, we introduce an efficient state truncation technique that connects our model to a general birth-and-death process with state-dependent rates. This simplifies the detailed balancing equations into a significantly smaller linear system of equations, facilitating the efficient calculation of steady-state distributions. These distributions are essential for evaluating general performance metrics, as acknowledged in Larson’s seminal work.
We emphasize that, akin to the rationale in Larson’s original paper (Larson 1974), no exact analytical expression exists for the hyperlattice model regarding travel time measures for an infinite-line capacity system. This challenge arises from the server’s position not being chosen based on the received region probability distribution of the call during system saturation periods when the server is dispatched consecutively. Consequently, in practice, we must assume that queueing delay is the primary factor influencing the server’s response time for each call.
Additionally, we highlight that our framework concentrates on a random policy controlled by a set of state-dependent matrices, and it can be extended to more intricate decision-making situations. This generalization would empower us to develop and optimize police dispatching policies that take into account queueing dynamics for more realistic systems. This consideration is particularly relevant since conventional First-Come-First-Served (FCFS) approaches may not be optimally efficient, leaving room for enhancement and optimization. Lastly, although our paper focuses on police districting, the generalized hypercube queueing model can also be applied to model mobile servers in a broader range of applications.
An important consideration emphasized in the literature is the explicit modeling of travel times in queueing systems, as region-dependent service rates and spatial dynamics critically influence queue behavior and dispatch effectiveness (e.g., Kanoria (2021), Besbes et al. (2022)). In our hyperlattice queueing model, travel times are treated as exogenous — consistent with Larson’s hypercube formulation — to maintain analytical and computational tractability. Nevertheless, incorporating travel-time dynamics directly into the Markov chain would improve accuracy for spatially sensitive dispatch policies. As an intermediate step, our supplementary experiments in Section EC2 incorporate heterogeneous service rates that reflect server-specific travel times, thereby validating the robustness of the truncated hyperlattice approximation. Extending the model to endogenize travel-time dynamics remains a promising direction for future research.